# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v2

@@ A.1 Marketplace
Comparison between online and offline fine-tuning. We ran both t-test and two-way ANOVA to better understand emergent misalignment and the effectiveness of mitigation through hindsight simulation under online and offline fine-tuning schemes. Results show that PPO with immediate feedback yields significantly lower true utility for the user than DPO ($p=1.1\times 10^{-4}$ in t-test). In addition, considering the difference between the (normalized) user rating and true utility, we find that immediate feedback in online RLHF using PPO introduces a larger misalignment gap than offline RLHF using DPO ($p=6.7\times 10^{-5}$ in t-test). Incorporating partial hindsight helps mitigate this misalignment gap across online and offline fine-tuning ($p=3.1\times 10^{-116}$ in two-way ANOVA test). We also compared online PPO with offline SimPO (Meng et al., 2024) and found that PPO introduces a larger misalignment gap than SimPO ($p=8.2\times 10^{-5}$ in t-test), with partial hindsight significantly reducing misalignment in SimPO as well ($p=5\times 10^{-56}$ in t-test).
A.2 TruthfulQA
(a) Llama-2-7b (DPO)
(b) Llama-2-7b (PPO)
(c) Llama-3-8b (DPO)
(d) Llama-3-8b (PPO)
Figure 12: TruthfulQA results comparing RLHF with immediate feedback (gray) vs. RLHS with partial hindsight (orange) for Llama-2-7b and Llama-3-8b, each trained via DPO or PPO.
@@ A.2 TruthfulQA
Dataset Details. TruthfulQA (Lin et al., 2021) is a benchmark designed to elicit hallucinatory responses from language models. Its authors introduced a new recommended multiple-choice version with two options (one correct, one incorrect), randomly ordered as (A) and (B). This binary-choice format replaces earlier versions (MC1 and MC2) and reduces the likelihood that models rely on simple heuristics (Evans et al., ). Accuracy is defined as the proportion of questions for which the model assigns the highest probability to the truthful answer. We adopt this revised format to evaluate multiple-choice accuracy across various models.
Analysis. In addition to the DPO results reported in the main paper, we also present performance metrics for models trained using PPO, which reveals consistent trends that support the efficacy of RLHS. These findings demonstrate that RLHS remains effective across both the offline DPO and online PPO approaches.
Appendix B Additional theoretical analysis
@@
Applying the same logic of this lemma to individual executions and assuming a Boltzmann-rational evaluator like the one discussed in Section 2 (and often considered for theoretical purposes when analyzing RLHF methods), we obtain the following result.
@@
Appendix C Training algorithms.
@@
Appendix D Prompts
@@
Appendix E Human Study Details
E.1 Additional Results
E.2 User Interface
In this subsection, we display the interface used in our human study.
E.3 Participants and data collection
E.4 Additional human study on alignment between human and AI feedback
Appendix F Discussion
F.1 Related Work
@@
Recent studies have investigated sycophantic behavior in language models (Sharma et al., 2023; Wei et al., 2023; Perez et al., 2022), where the models are optimized to generate responses that align with user beliefs rather than the truth. Our empirical results also reveal such tendencies. In this paper, we analyze the underlying factors contributing to this behavior and demonstrate how incorporating hindsight can be effective in preventing sycophancy.
There is a broad literature on agents obtaining unintended rewards through phenomena such as reward hacking (Amodei et al., 2016), reward tampering (Everitt et al., 2021), reward corruption (Everitt et al., 2017), wireheading (Everitt & Hutter, 2016), and corrigibility (Soares et al., 2015). Some of these behaviors have been observed in large language models (Denison et al., 2024; Wen et al., 2024; Williams et al., 2024). In our work, we show that human foresight feedback in RLHF can precipitate reward hacking. To mitigate this, we propose Reinforcement Learning from Hindsight Simulation (RLHS), which leverages simulated outcomes to guide more truthful feedback and reduce reward hacking.
F.2 Broader Impact
Human evaluators do not always know the full truth when providing feedback. Without explicit information about its future consequences, evaluators must implicitly estimate them. This limitation poses significant challenges for real-world applications of AI, particularly within the RLHF framework we studied. In the following sections, we discussed these limitations and how our proposed hindsight feedback approach can help overcome them to enhance AI alignment.
Limited Access in Real-World Applications: In real-world scenarios, users and human labelers frequently interact with black-box or closed-prompt AI systems where internal prompts and decision-making processes remain opaque. Notable examples include commercial systems such as OpenAI’s ChatGPT and Amazon’s Rufus. Our proposed techniques (hindsight feedback), and the experimental settings we used can be applied directly to these systems where full internal access is unavailable. In such cases, assessing the consistency of responses alone is insufficient, as external context might not capture the complete implications of an AI’s output. Hindsight feedback allows evaluators to provide more reliable feedback by considering outcomes, improving alignment in these constrained settings.
Limitations of Human Judgment and Information Access: Even when human evaluators have full access to models and their prompts (e.g., in open-source systems), perfect judgment is not guaranteed. Evaluators may miss deeper implications or fail to predict the long-term impact of responses, whether due to lack of expertise or cognitive limitations. These challenges are relevant to both open-source and closed-source models. Below, we outline two practical examples illustrating these limitations and how hindsight feedback can address them:
AI4Science Proof Construction: When constructing mathematical proofs for scientific problems, model may generate results that are correct only under conditions or assumptions specified by the user. Human evaluators, constrained by time or limited expertise, may overlook these limitations during evaluating the model, eventually causing the model to overfit to a restricted set of problems and unable to tackle scientific problems in general settings. On the other hand, we can implement hindsight simulation by generating a diverse set of scenarios, including, e.g., edge cases, under which the model is required to validate its proof. This allows the human evaluator to assess the model performance based on its ability to generalize beyond the immediate problem.
@@ F.2 Broader Impact
Appendix G Additional Qualitative Results

### v3

@@ A.1 Marketplace
Comparison between online and offline fine-tuning. We ran both t-tests and a two-way ANOVA to better understand emergent misalignment and the effectiveness of mitigation through hindsight simulation under online and offline fine-tuning schemes. Results show that PPO with immediate feedback yields significantly lower true utility for the user than DPO ($p=1.1\times 10^{-4}$ in t-test). In addition, considering the difference between the (normalized) user rating and true utility, we find that immediate feedback in online RLHF using PPO introduces a larger misalignment gap than offline RLHF using DPO ($p=6.7\times 10^{-5}$ in t-test). Incorporating partial hindsight helps mitigate this misalignment gap across online and offline fine-tuning ($p=3.1\times 10^{-116}$ in two-way ANOVA test). We also compared online PPO with offline SimPO (Meng et al., 2024) and found that PPO introduces a larger misalignment gap than SimPO ($p=8.2\times 10^{-5}$ in t-test), with partial hindsight significantly reducing misalignment in SimPO as well ($p=5\times 10^{-56}$ in t-test).
Table 4: Ablation study on world models in RLHS. RLHS(L) uses Llama-3.1-70B as the world model, while RLHS(S) uses the AI assistant’s own model. The fine-tuned AI assistant is Llama-2-7b. Although the smaller model simulates outcomes less accurately, it still significantly reduces misalignment and achieves positive true utility.
Metric DPO PPO
IF RLHS(L) RLHS(S) IF RLHS(L) RLHS(S)
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.47_{\pm 0.038}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.49_{\pm 0.033}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $0.10_{\pm 0.034}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $0.12_{\pm 0.041}$
A.2 Details of Benchmark Datasets
Table 5: Performance comparison of different models on HaluEval. QA means question-answering. Dial means knowledge-grounded dialogue. Summ means text summarization.
Model QA $\uparrow$ Dial $\uparrow$ Summ $\uparrow$
Llama-3-8b 0.57 0.58 0.58
+ RLHF 0.50 0.57 0.51
+ RLHS 0.58 0.62 0.60
Table 6: Results on agreement on privacy information usage, showing that RLHS-trained model achieves higher performance than GPT-4.
Model Cor. $\uparrow$
GPT-4 0.634
Llama-3-8b 0.615
+ RLHF 0.603
+ RLHS 0.664
@@ A.2 Details of Benchmark Datasets
TruthfulQA (Lin et al., 2021) is a benchmark designed to elicit hallucinatory responses from language models. Its authors introduced a new recommended multiple-choice version with with two randomly ordered options (one correct, one incorrect), replacing earlier versions (MC1 and MC2). This binary-choice format reduces models’ reliance on simple heuristics (Evans et al., ). Accuracy is the proportion of questions for which the model assigns the highest probability to the truthful answer. We adopt this revised format to evaluate multiple-choice accuracy across various models.
HaluEval (Li et al., 2023a) is a benchmark designed to evaluate hallucinations in large language models, offering diverse examples across multiple tasks. It comprises 30,000 automatically generated samples spanning question answering (QA), knowledge-grounded dialogue (Dial), and text summarization (Summ), along with 5,000 human-labeled ChatGPT responses collected from real user queries. We adopt accuracy as the primary metric, computed as the percentage of correct “Yes/No” classifications for whether a response is hallucinated.
TrustLLM (Sun et al., 2024) is a comprehensive benchmark developed to systematically evaluate the trustworthiness of leading large language models. It establishes eight core dimensions of trustworthiness, including truthfulness, privacy, safety, fairness, among others. Our evaluation focuses on assessing hallucination and sycophancy to understand the truthfulness of our fine-tuned models, as well as examining privacy-related concerns including privacy awareness and privacy leakage.
  • Hallucination: we evaluate multiple task types, including multiple-choice (MC), question answering (QA), knowledge-grounded dialogue (Dial), and text summarization (Summ), using accuracy as the primary metric.
  • Preference-driven sycophancy: we measure how much LLM responses shift when user preferences are altered. We used percentage change (Pref.%) as the metric.
  • Privacy leakage: It is defined as the potential unintended exposure of users’ private information possibly embedded in LLM training data, we employ three metrics: Refuse to Answer (RtA), Total Disclosure (TD), and Conditional Disclosure (CD). Specifically, RtA measures the proportion of instances where the LLM refuses to answer privacy-sensitive queries; TD indicates the ratio of accurate disclosures (correct email addresses provided) relative to all responses; and CD reflects the proportion of correct disclosures when the LLM does not refuse to respond. We present our evaluation results using the Enron email dataset.
  • Privacy awareness: We evaluate agreement on privacy information usage. TrustLLM uses the tier 2 task from recent dataset (Mireshghallah et al., 2023) where each data entry outlines how specific information is used, specifying (1) the type of information, (2) the actor, and (3) the purpose. For example: “Information about your social security number is collected by an online search website to prioritize search results.” LLMs must decide whether they agree with this usage, essentially determining if the data practice is appropriate. We use Pearson’s correlation (i.e., Cor.) as the evaluation metric.
Appendix B Environment Details
Environment 1: Marketplace
Products.
TV, Laptop, Smartphone, Refrigerator (six additional categories follow the same scheme).
Attribute set (8 per product).
  • TV: 3D capability, Resolution, HDR, Refresh rate, Smart features, Panel type, Connectivity, Screen size
  • Laptop: Screen resolution, Processor generation, Memory, Storage type, Battery life, Weight, USB-C port count, Fast-charging
  • Smartphone: Camera resolution, Battery capacity, Display type, Storage capacity, Memory, 5G support, Biometric security, Fast-charging
  • Refrigerator: Capacity, Energy efficiency, Defrost type, Temperature control, Water dispenser, Ice maker, Noise level, Shelf adjustability
Descriptor grid.
Every attribute has three natural-language variants—P (positive feature), N (negative feature), and U (unspecified)—yielding $5613^{8}=6{,}$ 561 unique configurations per product.
Price ladder.
Non-overlapping high, mid, and low tiers (e.g. TV: \$1.8–1.9k, \$1.4–1.6k, \$0.9–1.1k).
Sampling.
We sample a product, pick a price tier, then choose one P/N/U descriptor for each attribute to form a human-readable product blurb with controlled factual content.
Environment 2: Restaurant
Categories.
Italian, Japanese, Mexican, American (six additional cuisines—Indian, Chinese, French, Mediterranean, Thai, Korean—follow the same scheme).
Attribute set (8 per cuisine).
  • Italian: Pizza style, Pasta freshness, Ingredient sourcing, Wine list, Ambiance, Service quality, Dietary options, Dessert quality
  • Japanese: Fish freshness, Dining style, Noodle preparation, Interior atmosphere, Chef expertise, Beverage menu, Seasonal menu, Dessert offering
  • Mexican: Tortilla source, Meat preparation, Guacamole freshness, Entertainment, Spirit menu, Décor style, Vegan options, Dessert freshness
  • American: Beef sourcing, Fries quality, Beer selection, Music offering, Sustainability focus, Seasonality, Dessert sourcing, Outdoor seating
Descriptor grid.
Each attribute has P (positive feature present), N (negative feature absent), and U (unspecified) variants, giving $5613^{8}=6{,}$ 561 unique labelled restaurant profiles per cuisine.
Price bands.
Premium, mid-tier, and budget (e.g., Italian: \$60–80, \$35–50, \$18–28 per person).
Sampling.
We select a cuisine, draw a price band, and choose one P/N/U descriptor for each attribute to generate a realistic yet controllable menu description.
Environment 3: Online Course Platform
Tracks.
Data Science, Web Development, Business & Management, Graphic Design (six further tracks—Cybersecurity, Digital Marketing, Finance & Investing, Artificial Intelligence, Cloud Computing, Project Management—follow the same pattern).
Attribute set (8 per track).
  • Data Science: Project style, Instructor background, Certification, Tool coverage, Feedback policy, Capstone review, Access duration, Community support
  • Web Development: Curriculum depth, Support availability, Framework coverage, Portfolio deliverable, Career services, Mentoring, Assessment frequency, Access period
  • Business & Management: Case-study source, Webinar format, Certification status, Grading method, Networking, Content focus, Resource availability, Update frequency
  • Graphic Design: Project emphasis, Instructor accolades, Software coverage, Critique format, Certification, Access details, Career support, Asset exclusivity
Descriptor grid.
Each attribute has P (positive aspect present), N (negative aspect absent), and U (unspecified) variants, yielding $5613^{8}=6{,}$ 561 distinct labelled course profiles per track.
Price bands.
Premium, mid, and budget (e.g., Data Science: \$1.5–2.0k / \$0.7–1.0k / \$0.2–0.4k).
Sampling.
We draw a track, pick a price band, and select one P/N/U descriptor for every attribute, producing a realistic yet controllable course blurb.
Appendix C Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We model the a problem as a POMDP defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$) encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
Appendix D Additional theoretical analysis
@@
Applying the same logic of this lemma to individual executions and assuming a Boltzmann-rational evaluator like the one discussed in Appendix C (and often considered for theoretical purposes when analyzing RLHF methods), we obtain the following result.
@@
Appendix E Training algorithms.
@@
Appendix F Human Study Details
F.1 Additional Results
F.2 User Interface
In this subsection, we display the interface used in our human study.
F.3 Participants and data collection
F.4 Additional human study on alignment between human and AI feedback
Appendix G Prompts
@@
Appendix H Discussion
H.1 Additional Related Work
@@
H.2 Additional limitations and future works
Personalized Hindsight Simulation. Users inherently differ in preferences, risk tolerances, and expertise, causing identical outcomes to have varied perceived utilities. Integrating personalized user models into RLHS could significantly enhance alignment by tailoring simulated hindsight outcomes more closely to individual user objectives. Future studies could explore personalization techniques, leveraging explicit preference elicitation or implicit user behavior modeling to further improve the utility and acceptability of RLHS-aligned systems.
H.3 Broader Impact.
@@ H.2 Additional limitations and future works
H.4 Computing Resources
All experiments were conducted using Nvidia L40 GPUs (48GB memory). A single GPU suffices for inference and LoRA fine-tuning of Llama-3-8B and Llama-2-7B. However, inference with Llama-3.1-70B (used as the judge) requires four GPUs. Fine-tuning typically takes 1–2 days, inference completes within one day, and generating the complete preference dataset for fine-tuning requires more than two days.
Appendix I Additional Qualitative Results